Status Report

Terra Weekly Mission Status Update 5-11 Jan 2001

By SpaceRef Editor
January 11, 2001
Filed under ,

Terra is currently configured to operate in the nominal mode. All five of its instruments are in full Science Mode.

Terra operations were relocated to the Back-up EOS Operations Center (BEOC) at NASA*s Goddard Space Flight Center for one full shift on January 10, 2001. This transition was necessitated by a power outage in the primary EOS Operations Center (EOC) that was required to upgrade and verify the back-up power systems (Un-interruptible Power System, or UPS) in the building. The transition to the BEOC and back to the primary EOC was smooth and fully successful, as was the execution of the power system work. Additional UPS reconfiguration work is planned for January 24, 2001.

The Terra satellite is entering a period of high beta angle. The temperatures across the satellite are being monitored closely since this represents the completion of its first yearly thermal cycle.

The Terra flight operations team supported the MISR field campaign over JPL on January 12. It also support the MODIS snow and ice validation campaign over New York on January 15.


Work on the Solar Array shunt anomaly and the corruption of the Solar Array temperature sensors is continuing. Additionally, a team has been formed to work on the ASTER SWIR temperature yellow limits issue.


There will be a Drag Makeup Maneuver on January 17.

SpaceRef staff editor.